OpenStudy (anonymous):

solve 2/9 + 4/9 2/9 2/3 8/9 1/3

OpenStudy (cwrw238):

when the denominator ( bottom bit) of two fractions is the same just add the top bits - the denominator remains the same

OpenStudy (anonymous):

2/9+4/9=6/9 6/9=2/3